# Transversal Structures in Graph Systems: A Survey

### v1

@@ 1 Introduction
How global parameters of a graph, such as its degree or chromatic number, can influence its local substructures? Is there some sufficiently large average degree or chromatic number that forces a given substructure to appear? Questions of this type are one of the most natural problems in graph theory, and there are many profound and interesting results. These are collectively referred to as extremal graph theory. It is a branch of combinatorial mathematics that studies the maximal or minimal properties of graphs under certain conditions.
Many important problems in extremal graph theory can be framed as subgraph containment problems, which ask for conditions on a graph $G$ that ensure it contains a copy of a general graph $H$. For example, Mantel’s theorem states that every $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle; Turán [132] generalized it by determining the maximum number of edges in a graph that forbids $K_{r}$, where $K_{r}$ denotes a complete graph on $r$ vertices; Dirac [39] proved that every $n$-vertex graph with minimum degree at least $n/2$ contains a Hamilton cycle. Such a graph is called a Dirac graph. For more results, the reader is referred to the book of Bollobás [15].
In this paper, we survey some recent developments on transversal problems in graph systems. The notion of transversal appears in diverse branches of mathematics. A transversal $X$ over a system =$\mathcal{F}=\{F_{1},\ldots,F_{m}\}$ of objects refers to an object that intersects each $F_{i}$ exactly once. Here, the object can be any mathematical object, such as sets, spaces, set systems, and matroids, etc. Several classical theorems, which guarantee the existence of certain objects, have been reiterated in the context of transversals, demonstrating the possibility and impossibility of obtaining such an object as a transversal over a certain system. To name a few, transversals are extensively studied for Carathéodory’s theorem [14], Helly’s theorem [75], Erdős–Ko–Rado theorem [7], and Rota’s basis conjecture [71, 118], etc.
Building on this trend, there have been numerous studies dedicated to generalizing classical results in extremal graph theory to transversal version. In addition to being interesting in their own right, they often provide a strengthening of the original results, which also reflects the robustness of graph properties. A transversal in a graph system is also referred to as a “rainbow subgraph” in much literature, and it was formally introduced in the recent work of Joos and Kim [74]. For a positive integer $n$, we write $[n]:=\{1,2,\dots,n\}$.
@@
For a given system $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of graphs/digraphs/hypergraphs with the same vertex set $V$, an $m$-edge graph/digraph/hypergraph $H$ on the vertex set $V$ is a $\mathcal{G}$-transversal (or a transversal inside $\mathcal{G}$) if there exists a bijection $\phi:E(H)\to[m]$ such that $e\in E(G_{\phi(e)})$ for all $e\in E(H).$
@@
Perhaps the most famous transversals are those related to Latin squares, which were studied by Euler. In 1782, Euler [48] considered a Latin square of order $n$, which is an $n\times n$ array filled with symbols $1,\ldots,n$, where each symbol appears exactly once in each row and column. A partial transversal of a Latin square is a set of cells such that every pair of cells shares no row, column or symbol, in particular a transversal is a partial transversal with $n$ cells. The famous Ryser–Brualdi–Stein conjecture [124, 22, 126] is one of the most important problems on transversals in Latin squares.
@@
Every Latin square of order $n$ has a partial transversal with $n-1$ cells, and a transversal with $n$ cells if $n$ is odd.
@@
There is some $n_{0}\in\mathbb{N}$ such that every Latin square of order $n\geq n_{0}$ contains a partial transversal with $n-1$ cells.
@@
Considering the rows and columns of the Latin square as a complete bipartite graph $K_{n,n}$, where each symbol represents a color and each cell represents a colored edge, the Latin square naturally corresponds to a properly edge-colored $K_{n,n}$. The edge-colored $K_{n,n}$ can be viewed as a set of graphs $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$, where each $G_{i}$ consists of edges with color $i$, then a rainbow matching of $\mathcal{G}$ corresponds to a transversal of the Latin square. We refer the reader to see the nice survey [133]. These studies have inspired research on the existence of transversal structures in a system of graphs. The central question in the study of transversals can be generalized and depicted as follows.
@@
Let $H$ be a graph/digraph/hypergraph with $m$ edges, $\mathcal{G}=\{G_{1},G_{2},\dots,G_{m}\}$ be a system of graphs/digraphs/hypergraphs with order $n$. Which property imposed on $\mathcal{G}$ will ensure a transversal copy of $H$?
Since we do not require the objects in the family $\mathcal{G}$ to be distinct, a necessary condition for a positive answer to the above problem is that every $n$-vertex graph/digraph/hypergraph satisfying the property contains a copy of $H$. In the following of this survey, it will be shown by some examples that this necessary condition is not sufficient in some cases.
@@ 2 Turán-type problem in graph systems
One of the central topics in extremal combinatorics is to determine the maximum number of edges of a graph on $n$ vertices that does not contain a copy of a given graph $H$ as a subgraph, such a maximum number is denoted by $\mathrm{ex}(n,H)$. The most notable result is Turán’s theorem [132], which states that $\mathrm{ex}(n,K_{r+1})=t_{r}(n)$, where $t_{r}(n)$ is the number of edges of a balanced complete $r$-partite graph on $n$ vertices (denoted by $T_{r}(n)$). Erdős, Stone and Simonovits [46, 47] found a surprising connection between $\mathrm{ex}(n,H)$ and the chromatic number $\chi(H)$ of $H$.
In addition to the Turán problem itself, there are also many variants involving colorings. For example, Diwan and Mubayi [40] considered a generalization of Turán’s theorem in a graph system containing two graphs. It is of course possible to consider similar question with more than two graphs. In this section, we survey results on the Turán-type problems in graph systems. In other words, we are to determine the “largest” $n$-vertex graph systems that do not contain a copy of $H$ as a rainbow subgraph.
@@
Interestingly, the constant $\frac{26-2\sqrt{7}}{81}$ is an irrational number and larger than $\frac{1}{4}$. Therefore, compared to Mantel’s theorem, the transversal triangle problem requires a stronger condition. This signifies the difference between the transversal setting and the classical setting, and renders the rainbow extremal number more intriguing. Aharoni, DeVos, de la Maza, Montejano and Šámal [3] proposed a potential interesting direction to proceed, that is,
@@
For what real numbers $\alpha_{1},\alpha_{2},\alpha_{3}>0$, every graph system $\{G_{1},G_{2},G_{3}\}$ satisfying $|E(G_{i})|>\alpha_{i}n^{2}$ must have a transversal triangle?
Falgas-Ravry, Markström and Räty [50, Theorem 1.13] solved this problem when $n$ is sufficiently large. Recall that Turán’s theorem [132] generalizes Mantel’s theorem by determining the maximum number of edges in $n$-vertex $K_{r}$-free ($r\geq 3$) graphs. Analogously, one may consider the following.
@@
In general cases, the problem becomes more complicated and there are only few known results related to the $k$-rainbow extremal number $\mathrm{ex}_{k}^{*}(n,H)$. As a corollary of Keevash, Saks, Sudakov and Verstraëte [83], we have $\mathrm{ex}_{k}^{*}(n,K_{3})\leq n^{2}/4+o(n^{2})$ when $k\geq 4$. Frankl [55] gave a new proof of this result. Babiński and Grzesik [12] determined the value of $\mathrm{ex}_{k}^{*}(n,P_{4})$ up to additive $o(n^{2})$ error term for every $k\geq 3$, where $P_{4}$ denotes the path on $4$ vertices.
@@
Except the rainbow Turán number of a graph, depending on the contexts, different measures for the “size” of graph systems have been introduced, for example, the sum $\sum_{i}|E(G_{i})|$ [83, 24], the product $\Pi_{i}|E(G_{i})|$ [55, 68] and other measures.
The $k$-color Turán number, denoted by $\mathrm{ex}_{k}(n,H)$, is the maximum value of $\sum_{i\in[k]}|E(G_{i})|$ among all graph systems $\mathcal{G}=\{G_{1},\ldots,G_{k}\}$ of order $n$ and without rainbow copy of $H$. We call such a graph system $\mathcal{G}$ satisfying $\sum_{i\in[k]}|E(G_{i})|=\mathrm{ex}_{k}(n,H)$ the $k$-color extremal graph system of $H$. To determine $\mathrm{ex}_{k}(n,H)$, there are two natural lower bounds that arise from the following constructions. One is that each $G_{i}$ is the $n$-vertex $H$-free graph with $\mathrm{ex}(n,H)$ edges; the other one is that there are $|E(H)|-1$ copies of the complete graph $K_{n}$ and all the other graphs are empty.
A graph is called $r$-color-critical if it has chromatic number $r$, and it has an edge whose removal reduces the chromatic number to $r-1.$ In 2004, Keevash, Saks, Sudakov and Verstraëte [83] proposed the following conjecture.
@@
Suppose $r\geq 3$ and $k\geq h$. Let $H$ be an $r$-color-critical graph with $h$ edges. Then, there exists an $n_{0}=n_{0}(H)>0$ such that for all $n\geq n_{0}$, the $k$-color extremal graph system of $H$ either consists of exactly $h-1$ nonempty graphs where each of them is a copy of $K_{n}$ or consists of $k$ graphs where all of them are identical copies of $T_{r-1}(n)$. In particular,
@@
There is also some research focusing on maximizing the product $\Pi_{i}|E(G_{i})|$ of the number of edges of a graph system without a rainbow copy of $H$. Frankl [55] conjectured that a graph system $\{G_{1},G_{2},G_{3}\}$ on a common vertex set of order $n$ with $\Pi_{i\in[3]}|E(G_{i})|>\lfloor{n^{2}/4}\rfloor^{3}$ contains a transversal triangle. This conjecture was disproved by Frankl, Győri, He, Lv, Salia, Tompkins, Varga and Zhu [68], and they further conjectured that:
@@
In fact, this conjecture is a direct consequence of Problem 2.2, which has been solved by Falgas-Ravry, Markström and Räty [50] when $n$ is sufficiently large.
@@ 3 Dirac-type problem in graph systems
In this section, we investigate the minimum degree condition that ensure the existence of transversal spanning structures in graph systems. When one is interested in embedding a spanning graph, it makes sense to consider minimum degree conditions rather than size conditions, as, for example, a graph could be very dense but contain an isolated vertex, and therefore not contain a copy of any given spanning graph with minimum degree at least $1$.
Given the difficulty of obtaining exact results in (rainbow) graph embedding problems, many results have focused on determining the minimum $d$ such that for all $\epsilon>0$ and sufficiently large integer $n$, each graph (system) of order $n$ with minimum degree at least $(d+\epsilon)n$ contains a (transversal) copy of $H$. Note that when we consider this problem in graph systems, $d$ is at least the value required to guarantee a copy of $H$ in a single graph, since all graphs in the system could be identical.
@@
Let $H$ be a graph with $m$ edges. Denote by $\delta_{H}$ (if it exists) the smallest real number $\delta$ such that for all $\alpha>0$ and sufficiently large $n$, every $n$-vertex graph $G$ with $\delta(G)\geq(\delta+\alpha)n$ contains a copy of $H$.
Denote by $\delta_{H}^{rb}$ (if it exists) the smallest real number $\delta$ such that for all $\alpha>0$ and sufficiently large $n$, every graph system $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of order $n$ with $\delta(\mathcal{G})\geq(\delta+\alpha)n$ contains a transversal copy of $H$.
For a given graph $H$, if the above two values exist, then it is clear that $\delta_{H}^{\mathrm{rb}}\geq\delta_{H}.$ Indeed, if $G$ is $H$-free, then the graph system $\mathcal{G}$ consisting of $|E(H)|$ copies of $G$ does not contain a transversal copy of $H$. However, Montgomery, Müyesser and Pehova [115] showed that $\delta_{H}^{rb}$ can be much larger than $\delta_{H}$: taking $H$ to be the disjoint union of copies of $K_{2,3}\cup C_{4}$, then $\delta_{H}^{rb}\geq 1/2$, whereas Kühn and Osthus [95] showed that $\delta_{H}\leq 4/9$.
On the other hand, there are many natural instances with $\delta_{H}^{\mathrm{rb}}=\delta_{H}$, for example, Hamilton cycle and perfect matching (see [74]). When this equality holds, we always say that the corresponding $H$ is color-blind.

### v2

@@ 1 Introduction
How do global parameters of a graph, such as its degree or chromatic number, influence its local substructures? Is there some sufficiently large average degree or chromatic number that forces a given substructure to appear? Questions of this type are one of the most natural problems in graph theory, and there are many profound and interesting results. These are collectively referred to as extremal graph theory. It is a branch of combinatorial mathematics that studies the maximal or minimal properties of graphs under certain conditions.
Many important problems in extremal graph theory can be framed as subgraph containment problems, which ask for conditions on a graph $G$ that ensure it contains a copy of a general graph $H$. For example, Mantel’s theorem states that every $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle; Turán [140] generalized it by determining the maximum number of edges in an $n$-vertex graph that forbids $K_{r}$, where $K_{r}$ denotes a complete graph on $r$ vertices; Dirac [41] proved that every $n$-vertex ($n\geq 3$) graph with minimum degree at least $n/2$ contains a Hamilton cycle. For more results, the reader is referred to the book of Bollobás [15].
In this paper, we survey some recent developments on transversal problems in graph systems. The notion of transversal appears in diverse branches of mathematics. A transversal $X$ over a system $\mathcal{F}=\{F_{1},\ldots,F_{m}\}$ of objects refers to an object that intersects each $F_{i}$ exactly once. Here, the object can be any mathematical object, such as sets, spaces, set systems, and matroids, etc. Several classical theorems, which guarantee the existence of certain objects, have been reiterated in the context of transversals, demonstrating the possibility and impossibility of obtaining such an object as a transversal over a certain system. To name a few, transversals are extensively studied for Carathéodory’s theorem [14], Helly’s theorem [81], Erdős–Ko–Rado theorem [7], and Rota’s basis conjecture [77, 125], etc.
Building on this trend, there have been numerous studies dedicated to generalizing classical results in extremal graph theory to transversal versions. In addition to being interesting in their own right, they often provide a strengthening of the original results, which also reflects the robustness of graph properties. A transversal in a graph system is also referred to as a “rainbow subgraph” in much literature, and it was formally introduced in the recent work of Joos and Kim [80]. For a positive integer $n$, we write $n]:=\{1,2,\dots,n\}$.
@@
For a given system $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of graphs/digraphs/hypergraphs with the same vertex set $V$, an $m$-edge graph/digraph/hypergraph $H$ with vertices in $V$ is a $\mathcal{G}$-transversal (or a transversal inside $\mathcal{G}$) if there exists a bijection $\phi:E(H)\to[m]$ such that $e\in E(G_{\phi(e)})$ for all $e\in E(H).$
@@
Perhaps the most famous transversals are those related to Latin squares, which were studied by Euler. In 1782, Euler [50] considered a Latin square of order $n$, which is an $n\times n$ array filled with symbols $1,\ldots,n$, where each symbol appears exactly once in each row and column, see Figure (a). A partial transversal of a Latin square is a set of cells such that every pair of cells shares no row, column or symbol, in particular a transversal is a partial transversal with $n$ cells. The famous Ryser–Brualdi–Stein conjecture [22, 131, 133] is one of the most important problems on transversals in Latin squares.
@@
Every Latin square of order $n$ has a partial transversal with n1n-1 cells, and a transversal with $n$ cells if $n$ is odd.
@@
There is some $n_{0}\in\mathbb{N}$ such that every Latin square of order $n\geq n_{0}$ contains a partial transversal with n1n-1 cells.
@@
Considering the rows and columns of the Latin square as a complete bipartite graph $K_{n,n}$, where each symbol represents a color and each cell represents a colored edge, the Latin square naturally corresponds to a properly edge-colored $K_{n,n}$, see Figure (b). The edge-colored $K_{n,n}$ can be viewed as a set of graphs $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$, where each $G_{i}$ consists of edges with color $i$, then a rainbow matching of $\mathcal{G}$ corresponds to a transversal of the Latin square. We refer the reader to see the nice survey [141]. These studies have inspired research on the existence of transversal structures in a system of graphs. The central question in the study of transversals can be generalized and depicted as follows.
@@
Let $H$ be a graph/digraph/hypergraph with $m$ edges, and let $\mathcal{G}=\{G_{1},G_{2},\dots,\linebreak G_{m}\}$ be a system of graphs/digraphs/hypergraphs. Which property imposed on $\mathcal{G}$ will ensure a transversal copy of $H$?
Since we do not require the objects in the family $\mathcal{G}$ to be distinct, a necessary condition for a positive answer to the above problem is that every $n$-vertex graph/digraph/hypergraph satisfying the property contains a copy of $H$. In this survey, it will be shown by some examples that this necessary condition is not sufficient in some cases.
@@ 2 Turán-type problem in graph systems
One of the central topics in extremal combinatorics is to determine the maximum number of edges of a graph on $n$ vertices that does not contain a copy of a given graph $H$ as a subgraph, such a maximum number is denoted by $\mathrm{ex}(n,H)$. The most notable result is Turán’s theorem [140], which states that $\mathrm{ex}(n,K_{r+1})=t_{r}(n)$, where $t_{r}(n)$ is the number of edges of a balanced complete $r$-partite graph on $n$ vertices (denoted by $T_{r}(n)$). Erdős, Stone and Simonovits [48, 49] found a connection between $\mathrm{ex}(n,H)$ and the chromatic number $\chi(H)$ of $H$.
In addition to the Turán problem itself, there are also many variants involving colorings. For example, Diwan and Mubayi [42] considered a generalization of Turán’s theorem in a graph system containing two graphs. It is of course possible to consider similar questions with more than two graphs. In this section, we survey results on the Turán-type problems in graph systems. In other words, we are to determine the “largest” $n$-vertex graph systems that do not contain a copy of $H$ as a rainbow subgraph.
@@
Interestingly, the constant $\frac{26-2\sqrt{7}}{81}$ is an irrational number and larger than $\frac{1}{4}$. Therefore, compared to Mantel’s theorem, the transversal triangle problem requires a stronger size condition. This signifies the difference between the transversal setting and the classical setting, and renders the rainbow extremal number more intriguing. Aharoni, DeVos, de la Maza, Montejano and Šámal [3] proposed a potential interesting direction to proceed.
@@
For what real numbers $\alpha_{1},\alpha_{2},\alpha_{3}>0$, is it true that every graph system $\{G_{1},G_{2},G_{3}\}$ satisfying $|E(G_{i})|>\alpha_{i}n^{2}$ must have a transversal triangle?
Falgas-Ravry, Markström and Räty [52, Theorem 1.13] solved this problem when $n$ is sufficiently large. Recall that Turán’s theorem [140] generalizes Mantel’s theorem by determining the maximum number of edges in $n$-vertex $K_{r}$-free ($r\geq 3$) graphs. Analogously, one may consider the following.
@@
In general cases, the problem becomes more complicated and there are only few known results related to the $k$-rainbow extremal number $\mathrm{ex}_{k}^{*}(n,H)$. As a corollary of Keevash, Saks, Sudakov and Verstraëte [89], we have $\mathrm{ex}_{k}^{*}(n,K_{3})\leq n^{2}/4+o(n^{2})$ when $k\geq 4$. Frankl [57] gave a new proof of this result. Babiński and Grzesik [12] determined the value of $\mathrm{ex}_{k}^{*}(n,P_{4})$ up to an additive $o(n^{2})$ error term for every $k\geq 3$, where $P_{4}$ denotes the path on $4$ vertices.
@@
Except for the rainbow Turán number of a graph, depending on the contexts, different measures for the “size” of graph systems have been introduced, for example, the sum $\sum_{i}|E(G_{i})|$ [25, 89], the product $\Pi_{i}|E(G_{i})|$ [57, 74] and other measures.
The $k$-color Turán number, denoted by $\mathrm{ex}_{k}(n,H)$, is the maximum value of $\sum_{i\in[k]}|E(G_{i})|$ among all graph systems $\mathcal{G}=\{G_{1},\ldots,G_{k}\}$ of order $n$ and without rainbow copy of $H$. We call such a graph system $\mathcal{G}$ satisfying $\sum_{i\in[k]}|E(G_{i})|=\mathrm{ex}_{k}(n,H)$ the $k$-color extremal graph system of $H$. To determine $\mathrm{ex}_{k}(n,H)$, there are two natural lower bounds that arise from the following constructions. One is that each $G_{i}$ is the $n$-vertex $H$-free graph with $\mathrm{ex}(n,H)$ edges; the other one is that there are |E(H)|1|E(H)|-1 copies of the complete graph $K_{n}$ and all the other graphs are empty.
A graph is called $r$-color-critical if it has chromatic number $r$, and it has an edge whose removal reduces the chromatic number to r1.r-1. In 2004, Keevash, Saks, Sudakov and Verstraëte [89] proposed the following conjecture.
@@
Suppose $r\geq 3$ and $k\geq h$. Let $H$ be an $r$-color-critical graph with $h$ edges. Then, there exists an $n_{0}=n_{0}(H)>0$ such that for all $n\geq n_{0}$, the $k$-color extremal graph system of $H$ either consists of exactly h1h-1 nonempty graphs where each of them is a copy of $K_{n}$ or consists of $k$ graphs where all of them are identical copies of $T_{r-1}(n)$. In particular,
@@
There is also some research focusing on maximizing the product $\Pi_{i}|E(G_{i})|$ of the number of edges of a graph system without a rainbow copy of $H$. Frankl [57] conjectured that a graph system $\{G_{1},G_{2},G_{3}\}$ on a common vertex set of order $n$ with $\Pi_{i\in[3]}|E(G_{i})|>\lfloor{n^{2}/4}\rfloor^{3}$ contains a transversal triangle. This conjecture was disproved by Frankl, Győri, He, Lv, Salia, Tompkins, Varga and Zhu [74], and they further provided the following conjecture.
@@
Recall that Falgas-Ravry, Markström and Räty [52] solved Problem 2.2 for sufficiently large $n$. As a direct consequence, Conjecture 2.7 holds when $n$ is sufficiently large.
@@ 3 Dirac-type problem in graph systems
In this section, we investigate the minimum degree condition that ensures the existence of transversal spanning structures in graph systems. When one is interested in embedding a spanning graph, it makes sense to consider minimum degree conditions rather than size conditions, as, for example, a graph could be very dense but contain an isolated vertex, and therefore not contain a copy of any given spanning graph with minimum degree at least $1$.
Given the difficulty of obtaining exact results in (rainbow) graph embedding problems, many results have focused on determining the minimum $d$ such that for all $\epsilon>0$ and sufficiently large integer $n$, each graph (system) of order $n$ with minimum degree at least $(d+\epsilon)n$ contains a (rainbow) copy of $H$. Note that when we consider this problem in graph systems, $d$ is at least the value required to guarantee a copy of $H$ in a single graph, since all graphs in the system could be identical.
@@
Let $\mathbf{F}$ be an infinite family of graphs. Denote by $\delta_{\mathbf{F}}$ (if it exists) the smallest real number $\delta$ such that for all $\alpha>0$ and for all but finitely many $F\in\mathbf{F}$ the following holds. Let $n=|V(F)|$ and $H$ be any $n$-vertex graph with $\delta(H)\geq(\delta+\alpha)n$. Then $H$ contains a copy of $F$.
Denote by $\delta_{\mathbf{F}}^{rb}$ (if it exists) the smallest real number $\delta$ such that for all $\alpha>0$ and for all but finitely many $F\in\mathbf{F}$ the following holds. Let $n=|V(F)|$ and $m=|E(F)|$, and let $\mathcal{H}=\{H_{1},\ldots,H_{m}\}$ be any graph system of order $n$ with $\delta(\mathcal{H})\geq(\delta+\alpha)n$. Then $\mathcal{H}$ contains a transversal copy of $F$.
If the above two values exist, then it is clear that $\delta_{\mathbf{F}}^{\mathrm{rb}}\geq\delta_{\mathbf{F}}.$ Moreover, Montgomery, Müyesser and Pehova [122] observed that $\delta_{\mathbf{F}}^{rb}$ can be much larger than $\delta_{\mathbf{F}}$. Set $\mathbf{F}=\{k\times(K_{2,3}\cup C_{4}):k\in\mathbb{N}\}$, where $k\times G$ denotes the graph obtained by taking $k$ vertex-disjoint copies of $G$. Kühn and Osthus [101] showed that $\delta_{\mathbf{F}}=4/9$. Consider the graph system $\mathcal{H}=\{H_{1},\ldots,H_{m}\}$ on $V$ obtained in the following way. Partition $V$ into two almost equal vertex subsets, say $A$ and $B$, and for each $i\in[m-1]$, assume that $H_{i}$ is the disjoint union of a clique on $A$ and a clique on $B$. Suppose that $H_{m}$ is a complete bipartite graph between $A$ and $B$. Observe that if $\mathcal{H}$ contains a transversal copy of some $F\in\mathbf{F}$, the edge of $K_{2,3}$ or $C_{4}$ that belongs to $H_{m}$ would be a bridge of $F$. However, neither $K_{2,3}$ nor $C_{4}$ contains a bridge. Hence, $\delta_{\mathbf{F}}^{rb}\geq\delta(\mathcal{H})\geq 1/2$.
On the other hand, there are many natural instances with $\delta_{\mathbf{F}}^{\mathrm{rb}}=\delta_{\mathbf{F}}$, for example, the family of $n$-cycles (see [80]). When this equality holds, we say that the corresponding $\mathbf{F}$ is color-blind.